Julia and Mary each improved their yards by planting hostas and ivy. They bought their supplies from the same store. Julia spent $93 on 10 hostas and 9 pots of ivy. Mary spent $80 on 8 hostas and 8 pots of ivy. Find the cost one hosta and the cost of one pot of I've.
Answer by stanbon(75887)

Julia and Mary each improved their yards by planting hostas and ivy. They bought their supplies from the same store. Julia spent $93 on 10 hostas and 9 pots of ivy. Mary spent $80 on 8 hostas and 8 pots of ivy. Find the cost one hosta and the cost of one pot of I've.
-------
Equations:
10h + 9p = 93 dollars
8h + 8p = 80 dollars
------
Modify:
Multiply thru top equation by 8
Multiply thru bottom equation by 10
-------
80h + 72p = 8*93
80h + 80p = 800
-----
Subtract and solve for "p":
8p = 56
p = $7.00 (price of one pot of ivy)
-----
Solve for "h":
8h + 8p = 80
h + p = 10
h = $3.00 (price of one hosta plant)
====================
Cheers,
Stan H.
====================